# Description of the structure of DG–algebras via characters on a groupoid

### v2

@@ 2.1 Inner derivations
Consider the element $a\in\pm G$. Let’s define the mapping $\chi^{a}:Hom(\Gamma)\rightarrow\mathbb{C}$ as follows (assuming that $a\neq b$):
$\chi^{a}(\phi):=\begin{cases}1,&\text{$\phi\in Hom(b,a)$}\\ -1,&\text{$\phi\in Hom(a,b)$}\\ 0,&\text{else.}\end{cases}$ (5)
It is clear that $\chi^{a}$ is trivial on endomorphisms.
@@
Map $\chi^{a}$ sets the inner derivations for $a\in\pm G$ by formula
$d_{a}:x\rightarrow[a,x]_{grad}$ , where
$a,x]_{grad}=ax-(-1)^{|x|}xa$.
@@ 2.2 Central derivations
We will now give an example of a derivation that is not inner. It will be demonstrated that such a derivation has an easy calculation form and is not trivial on endomorphisms. It will thus be established that they cannot be presented as a (perhaps formal) sum of inner derivations.
@@
We say that map $\tau:G\rightarrow\mathbb{C}$ – is graded group character, if it satisfies the following property:
@@
Let’s fix the graded group character $\tau$, and let $z\in Z(G)$ be an element of group center.
@@
Define map $d_{z}^{\tau}:\mathbb{C}[G]\rightarrow\mathbb{C}[G]$ on the generators $g\in\pm G$ as
$d_{z}^{\tau}:g\rightarrow\tau(g)gz.$
Continued by linearity for the whole algebra $\mathbb{C}[G]$ map $d_{z}^{\tau}$ we will call it as the central graded derivation.
@@
[…]
@@
On arbitrary elements of a group algebra, the graded Leibniz rule is fulfilled due to the linearity of the $d_{z}^{\tau}$. ∎
@@
Let $\chi_{\tau,z}$ – a character that corresponds to derivation $d_{z}^{\tau}$. By corollary 1 we have equality
@@
Therefore, the character is not equal to 0 only in the case of $t=z$:
@@
Denote $\phi=((-1)^{|g|}gz,g)$ and calculate its beginning and end
@@
Then in the case when $g\in A_{2n}$, we have: $s(\phi)=t(\phi)$ $\Rightarrow\chi_{\tau,z}$ non trivial on a loop $\phi$, and $d_{z}^{\tau}$ it is not an inner derivation. ∎
A direct check shows that inner and central derivations are subalgebras. For more details see [Aru20].
@@ 2.3 Quasi-inner derivations
As seen from formula (5) inner derivations are trivial on endomorphisms; however, the opposite is not true. We will call the class of such derivations quasi-inner.
@@
Define QInnDer($\mathbb{C}[G]$) derivation as follows:
$QInnDer(\mathbb{C}[G])=\left\{d\in Der\mid\forall a\in Obj(\Gamma),\forall\phi\in Hom(a,a)\Rightarrow\chi_{d}(\phi)=0\right\}.$ (6)
Quasi-inner derivations are clearly a sub-space, but the more significant statement is that they constitute an ideal in algebra of all derivations.
@@
QInnDer($\mathbb{C}[G]$) $\triangleleft$ Der($\mathbb{C}[G]$) is an ideal.
@@
Let’s proof that QInnDer($\mathbb{C}[G]$) is a subalgebra, i.e.
$d_{1}$, $d_{2}$ $\in$ QInnDer($\mathbb{C}[G]$) $\Rightarrow$ $\left[d_{1},d_{2}\right]\in QInnDer(\mathbb{C}[G])$.
character $\chi_{d_{1}}$ can be represented as:
@@
since $\chi_{d_{1}}$ is trivial on endomorphisms, where $\chi_{d_{1}}$ defined by formula (5). Similarly with a character $\chi_{d_{2}}$:
@@
$\{\chi_{d_{1}},\chi_{d_{2}}\}\in QInnDer(\mathbb{C}[G])$, to proof that, consider the value of the character on the loop $(uz,z),z\in Z_{G}(u)$
@@
Now pass directly to the proof of the theorem. Represent $\chi_{d_{0}}$ as
@@
Prove the theorem for $\chi^{a}$ and extend the result to the $\chi_{d_{0}}$, using the bilinearity of the commutator. Consider character $\{\chi_{d},\chi^{a}\}$. Proof that it is trivial on endomorphisms, i.e. $\forall b\in G$ and for $\forall z\in Z_{G}(b)$ is performed as $\{\chi_{d},\chi^{a}\}(bz,z)=0$. By Proposition 3.1:
@@
$\chi^{a}(h,z)\neq 0$ only in 2 cases: when $h=za$ and $h=az$. $\chi^{a}(bz,h)\neq 0$ only in 2 cases: when $h=bza^{-1}$ and $h=a^{-1}bz$. It means that
@@
It is easy to see that central derivations are not quasi-inner, for which it is sufficient to repeat the reasoning from proposition 2.2.
@@
Our next goal is to give the description of possible DG–algebra sctructure up to isomorphism. We follow the definitions from [FHT01]
@@
Differential graded algebra (DG–algebra) is a graded algebra $A\cong\bigoplus A_{i}$, equipped with derivation
$d:A\rightarrow A$,
which has a degree of $1$ (the case of a cochain complex) or 1-1 (the case of a chain complex) and satisfies following properties
@@
We will give the necessary and sufficient conditions from the point of view of characters on a groupoid that derivatiation sets the structure of a DG–algebra on a graded group algebra. We will describe the case of a cochain complex, since the reasoning is similar for a chain complex.
@@
The characters on the groupoid define the structures of the DG–algebra if and only if the following conditions are satisfied:
@@
The main idea is to prove that the first condition is equivalent to $d^{2}(g)=0$, and that the second condition is equivalent to $d(A^{i})\subseteq A^{i+1}$.
@@
Using proposition 2.2, we rewrite the properties of the DG–algebra in terms of characters.
@@
    Let $g\in A^{i}$, than $d(g)=\sum\limits_{h\in G}(h,g)h\in A^{i+1}\Leftrightarrow(h,g)=0$, if $h\notin A^{i+1}$.
    Note that if $u\in A^{i+1},v\in A^{i}$, than for $\phi=(u,v)$ we have: $s(\phi)=(-1)^{|v|}v^{-1}u\in A^{1},t(\phi)=uv^{-1}\in A^{1}$,this means that in the case of a DG algebra, the characters may not be equal to 0 only on morphisms between elements from $A^{1}$.
Now let us recall the definition of an isomorphism with respect to the structure of DG–algebra.
@@
An isomorphism $f$ between two DG–algebras $(A_{1},d_{1})$ and $(A_{2},d_{2})$, $f$: $A_{1}$ $\to$ $A_{2}$, is an algebra isomorphism such that following properties hold:
@@
    f holds the structure of graded algebra $f(A^{i}_{1})\subset A^{i}_{2}$.
@@
    for all $a\in A_{1}$ holds $f(d_{1}(a))=d_{2}(f(a))$.
Now we will describe the structures of DG–algebras up to isomorphism from definition (3.2).
Let $X_{i},i=1,2$ be a matrix (if $G$ is an infinite group, that this matrix is also infinite) in following sense
$X:G\times G\to\mathbb{C},$
$(x,h)\mapsto\chi(x,h).$
@@
Two DG–algebras $(\mathbb{C}[G],d_{1})$ and $(\mathbb{C}[G],d_{2})$ are isomorphic iff there is a matrix C, such that
@@
    The mapping f is $Aut(\mathbb{C}[G])$.
In other words that means that sets of constants $c_{h}^{g}$, gives $f(g)=\sum\limits_{h\in G}c_{h}^{g}h\in Aut(\mathbb{C}[G])$ and
@@
Suppose that we have $f:(\mathbb{C}[G],d_{1})\rightarrow(\mathbb{C}[G],d_{2})$ — an isomorphism of DG–algebras given on generators by the formula $f(g)=\sum\limits_{h\in G}c_{h}^{g}h$. By assumption we have:
@@
$f(d_{1}(g))=f(\sum\limits_{h\in G}\chi_{1}(h,g)h)=\sum\limits_{h\in G}\chi_{1}(h,g)f(h)=\sum\limits_{h\in G}\chi_{1}(h,g)(\sum\limits_{x\in G}c_{x}^{h}x)=(\sum\limits_{h\in G}\chi_{1}(h,g)c_{x}^{h})x$;
$d_{2}(f(g))=d_{2}(\sum\limits_{h\in G}c_{h}^{g}h)=\sum\limits_{h\in G}c_{h}^{g}d_{2}(h)=\sum\limits_{h\in G}c_{h}^{g}(\sum\limits_{x\in G}\chi_{2}(x,h)x)=(\sum\limits_{h\in G}c_{h}^{g}\chi_{2}(x,h))x$.
From where we get our equality (8). To prove the converse, just define $f(g)=\sum\limits_{h\in G}c_{h}^{g}h$ and for the same reasons, we get the necessary equalities. ∎
@@
In the case when $f\in Aut(G)$ equality (8) takes the form
@@
In group algebra $\mathbb{C}[G]\cong\bigoplus A^{i}$, such that $A^{1}\neq 0$ and $\exists a\in Z(G)$ and $a\in A^{1}$, the inner derivation generated by the element $a$ according to the formula from proposition 4.1 defines the structure of the DG–algebra on $\mathbb{C}[G]$. Note that identity element of group lies in $A^{0}$.
@@
It must be shown that $d^{2}=0$ is fulfilled for such derivations. We know from Theorem (3.1) that this condition is equivalent to
$\sum\limits_{h\in G}\chi(h,g)\chi(x,h)=0$, $\forall g,x\in G.$
Consider the product $\chi(h,g)\chi(x,h)$, by definition of inner derivations, this product is not equal to 0 if and only if $(h,g)$ and $(x,h)$ $\in Hom(a,-a)$ or Hom(a,a)Hom(-a,a).
Let $s(h,g)=s(x,h)=a$ and t(h,g)=t(x,h)=at(h,g)=t(x,h)=-a, then $(-1)^{|g|}g^{-1}h=(-1)^{|h|}h^{-1}x$ and $hg^{-1}=xh^{-1}$, but it is impossible because $a\in A^{1}\Rightarrow|h|\neq|g|.$
Let $s(h,g)=t(x,h)=a$ and t(h,g)=s(x,h)=at(h,g)=s(x,h)=-a, then $(-1)^{|g|}g^{-1}h=xh^{-1}$ and $hg^{-1}=(-1)^{|h|}h^{-1}x,$ but this is similarly impossible because $|h|\neq|g|.$
@@
Let $d^{\tau}_{z}(g)=\tau(g)gz$ - the central derivation given by the element $z\in A^{1}$. It defines the structure of a DG–algebra if and only if $\tau(g)\tau(gz)=0,\forall g\in G$.
@@
Sets the structure of the DG–algebra on $\mathbb{C}[G]$.
@@
To prove, it is necessary to check only the fulfillment of the graduated Leibniz rule. And to do this, we need to make sure that the given $\tau$ is really a group character and take advantage of Proposition (3.1). It is not difficult to do this by simply going through all possible options. ∎
@@
In group graded algebra $\mathbb{C}[Z]\cong\bigoplus\mathbb{C}[x^{k}]$, where $\mathbb{Z}=\left\langle{x}\right\rangle$, it is possible to define the structure of the DG algebra using central derivation $d^{\tau}_{x}$, where
@@
Consider the Heisenberg integer group of 3×3 upper triangular matrices
@@
It is well known that any element of this group can be represented as $y^{b}z^{c}x^{a}$. Then $f(y^{b}z^{c}x^{a})=c\in\mathbb{Z}$ - an epimorphism into a group of integers. From here we get the graduation $\mathbb{C}[H_{3}]\cong\bigoplus_{k\in\mathbb{Z}}\mathbb{C}[z^{k}\left\langle{x,y}\right\rangle]$. And the central differentiation $d^{\tau}_{z}$, constructed similarly to Example 5.2, gives the structure of the DG–algebra on $\mathbb{C}[H_{3}]$.

### v3

@@ 2.1 Inner derivations
Consider an element $a\in\pm G$. Define the mapping $\chi^{a}:\operatorname{Hom}(\Gamma)\rightarrow\mathbb{C}$ as follows (assuming $a\neq b$):
$\chi^{a}(\phi):=\begin{cases}1,&\text{$\phi\in\operatorname{Hom}(b,a)$}\\ -1,&\text{$\phi\in\operatorname{Hom}(a,b)$}\\ 0,&\text{otherwise.}\end{cases}$ (5)
Clearly $\chi^{a}$ is trivial on endomorphisms.
@@
The map $\chi^{a}$ defines an inner derivations for $a\in\pm G$ by the formula
$d_{a}:x\mapsto[a,x]_{grad}=ax-(-1)^{|x|}xa$.
@@ 2.2 Central derivations
We will now give an example of a derivation that is not inner. We will show that it admits an simple explicit form and is not trivial on endomorphisms. That entails that it cannot be expressed as a (perhaps formal) sum of inner derivations.
@@
That the map $\tau:G\rightarrow\mathbb{C}$ is a graded group character if it satisfies the following property:
@@
Let’s fix a graded group character $\tau$, and let $z\in Z(G)$ be central element of the group.
@@
Define the map $d_{z}^{\tau}:\mathbb{C}[G]\rightarrow\mathbb{C}[G]$ on the generators $g\in\pm G$ as
$d_{z}^{\tau}:g\rightarrow\tau(g)gz$
and extend it by linearity to the whole algebra $\mathbb{C}[G]$. We will call it a central graded derivation.
@@
It is enough to check the graded Leibniz rule on arbitrary generators $u,v\in G$:
@@
On arbitrary elements of a group algebra, the graded Leibniz rule is satisfied due to the linearity of $d_{z}^{\tau}$. ∎
@@
Let $\chi_{\tau,z}$ be a character that corresponds to the derivation $d_{z}^{\tau}$. By corollary 1 we have the equality
@@
Therefore, this character is not equal to 0 only in the case $t=z$:
@@
Define the morphism $\phi=((-1)^{|g|}gz,g)$ and calculate its beginning and end
@@
Then in the case when $g\in A_{2n}$, we have: $s(\phi)=t(\phi)$ $\Rightarrow\chi_{\tau,z}$ is non–trivial on a loop $\phi$, thus $d_{z}^{\tau}$ it is not an inner derivation. ∎
A direct check shows that inner and central derivations constitute subalgebras. For more details see [Aru20].
@@ 2.3 Quasi-inner derivations
As seen from the formula (5), inner derivations are trivial on endomorphisms; however, the converse is not true. We will call such derivations quasi-inner.
@@
Define $\operatorname{QInnDer}(\mathbb{C}[G])$ as follows:
$\operatorname{QInnDer}(\mathbb{C}[G])=\left\{d\in\operatorname{Der}\mid\forall a\in\operatorname{Obj}(\Gamma),\forall\phi\in\operatorname{Hom}(a,a):\chi_{d}(\phi)=0\right\}.$ (6)
Quasi-inner derivations are clearly a subspace, but the more significant statement is that they constitute an ideal in the algebra of all derivations.
@@
$\operatorname{QInnDer}(\mathbb{C}[G])$ $\triangleleft$ $\operatorname{Der}(\mathbb{C}[G])$ is an ideal.
@@
Let us prove that $\operatorname{QInnDer}(\mathbb{C}[G])$ is a subalgebra, i.e.
$d_{1}$, $d_{2}$ $\in$ $\operatorname{QInnDer}(\mathbb{C}[G]$) $\Rightarrow$ $\left[d_{1},d_{2}\right]\in\operatorname{QInnDer}(\mathbb{C}[G])$.
The character $\chi_{d_{1}}$ can be represented as
@@
where $\chi_{d_{1}}$ defined by formula (5), since $\chi^{a}$ is trivial on endomorphisms. Similarly for the character $\chi_{d_{2}}$:
@@
thus $\{\chi_{d_{1}},\chi_{d_{2}}\}\in\operatorname{QInnDer}(\mathbb{C}[G])$. To prove that, consider the value of the character on the loop $(uz,z),z\in Z_{G}(u)$:
@@
We will now turn to the proof of the theorem. Represent $\chi_{d_{0}}$ as
@@
It is sufficient to prove the theorem for $\chi^{a}$ and extend the result to the $\chi_{d_{0}}$ using bilinearity of the commutator. Consider the character $\{\chi_{d},\chi^{a}\}$. We need to prove that it is trivial on endomorphisms, i.e. that for all $b\in G$ and for $z\in Z_{G}(b)$ it satisfies $\{\chi_{d},\chi^{a}\}(bz,z)=0$. By Proposition 3.1:
@@
Here $\chi^{a}(h,z)\neq 0$ only in 2 cases: when $h=za$ and when $h=az$; and $\chi^{a}(bz,h)\neq 0$ only in 2 cases: when $h=bza^{-1}$ and when $h=a^{-1}bz$. This means that
@@
It is easy to see that central derivations are not quasi-inner by repeating the reasoning from proposition 2.3.
@@
Our next goal is to give a description of the possible DG–algebra structures up to isomorphism. We follow the definitions of [FHT01].
@@
Let $A\cong\bigoplus A_{i}$ be a $\mathbb{Z}$–graded algebra equipped with a derivation
$d:A\rightarrow A$
of degree $1$ (the case of a cochain complex) or 1-1 (the case of a chain complex) and satisfying the following conditions:
@@
We say that $(A,d)$ is a differential graded algebra or DG–algebra.
We will give the necessary and sufficient conditions from the point of view of groupoid characters for a derivation to define a DG–algebra structure on a graded group algebra. We will describe the cochain complex case, since the reasoning in the chain complex is similar.
@@
A character $\chi$ on the groupoid $\Gamma$ defines a DG–algebra structure if and only if the following conditions are satisfied:
@@
The main idea is to prove that the first condition is equivalent to $d^{2}(g)=0$ and the second one is equivalent to $d(A^{i})\subseteq A^{i+1}$.
@@
Using proposition 2.2, we rewrite the defining properties of a DG–algebra in terms of characters.
@@
    Let $g\in A^{i}$, then $d(g)=\sum\limits_{h\in G}\chi(h,g)h\in A^{i+1}\Leftrightarrow\chi(h,g)=0$ when $h\notin A^{i+1}$.
    Note that if $u\in A^{i+1},v\in A^{i}$, then for $\phi=(u,v)$ we have: $s(\phi)=(-1)^{|v|}v^{-1}u\in A^{1}$, $t(\phi)=uv^{-1}\in A^{1}$,which means that in the case of a DG–algebra, the characters may not be equal to 0 only on morphisms between elements from $A^{1}$.
Now let us recall the definition of an isomorphism of DG–algebras.
@@
An isomorphism $f$: $A_{1}$ $\to$ $A_{2}$ between two DG–algebras $(A_{1},d_{1})$ and $(A_{2},d_{2})$ is an algebra isomorphism such that following properties hold:
@@
    f respects the grading, $f(A^{i}_{1})\subset A^{i}_{2}$;
@@
    for all $a\in A_{1}$, $f(d_{1}(a))=d_{2}(f(a))$.
Now we will describe the structures of DG–algebras up to isomorphism per definition (3.2).
Let $X_{i},i=1,2$ be two matrices (infinity ones, if $G$ is an infinite group) in following sense:
$X_{i}:G\times G\to\mathbb{C},$
$(x,h)\mapsto\chi_{i}(x,h).$
@@
Two DG–algebras $(\mathbb{C}[G],d_{1})$ and $(\mathbb{C}[G],d_{2})$ are isomorphic iff there is a matrix $C$ such that
@@
    the mapping f is in $\operatorname{Aut}(\mathbb{C}[G])$.
In other words, that means that sets of constants $c_{h}^{g}$, gives $f(g)=\sum\limits_{h\in G}c_{h}^{g}h\in\operatorname{Aut}(\mathbb{C}[G])$ and
@@
Suppose that we have an isomorphism of DG–algebras $f:(\mathbb{C}[G],d_{1})\rightarrow(\mathbb{C}[G],d_{2})$ ,given on generators by the formula $f(g)=\sum\limits_{h\in G}c_{h}^{g}h$. By definition, we have
@@
$f(d_{1}(g))=f\Big{(}\sum\limits_{h\in G}\chi_{1}(h,g)h\Big{)}=\sum\limits_{h\in G}\chi_{1}(h,g)f(h)=\sum\limits_{h\in G}\chi_{1}(h,g)\big{(}\sum\limits_{x\in G}c_{x}^{h}x\Big{)}=\Big{(}\sum\limits_{h\in G}\chi_{1}(h,g)c_{x}^{h}\Big{)}x$;
$d_{2}(f(g))=d_{2}\Big{(}\sum\limits_{h\in G}c_{h}^{g}h\Big{)}=\sum\limits_{h\in G}c_{h}^{g}d_{2}(h)=\sum\limits_{h\in G}c_{h}^{g}(\sum\limits_{x\in G}\chi_{2}(x,h)x)=\Big{(}\sum\limits_{h\in G}c_{h}^{g}\chi_{2}(x,h)\Big{)}x$.
Whence we get our equality (8). To prove the converse, just define $f(g)=\sum\limits_{h\in G}c_{h}^{g}h$ and for the same reasons, get the necessary equalities. ∎
@@
In the case when $f\in\operatorname{Aut(G)}$ the equality (8) takes the form
@@
In a group algebra $\mathbb{C}[G]\cong\bigoplus A^{i}$ such that $A^{1}\neq 0$, and given $a\in Z(G)\cap A^{1}$, the inner derivation generated by the element $a$ according to the formula from proposition 4.1 defines a DG–algebra structure on $\mathbb{C}[G]$. Note that the identity element of the group lies in $A^{0}$.
@@
We have to show that $d^{2}=0$ for such a derivation. We know from Theorem (3.1) that this condition is equivalent to
$\sum\limits_{h\in G}\chi(h,g)\chi(x,h)=0$; $\forall g,x\in G.$
Consider the product $\chi(h,g)\chi(x,h)$. By the definition of an inner derivation, this product is not equal to 0 if and only if $(h,g)$ and $(x,h)$ $\in Hom(a,-a)$ or Hom(a,a)Hom(-a,a).
Let $s(h,g)=s(x,h)=a$ and t(h,g)=t(x,h)=at(h,g)=t(x,h)=-a, then $(-1)^{|g|}g^{-1}h=(-1)^{|h|}h^{-1}x$ and $hg^{-1}=xh^{-1}$; but this is impossible because $a\in A^{1}\Rightarrow|h|\neq|g|.$
Let $s(h,g)=t(x,h)=a$ and t(h,g)=s(x,h)=at(h,g)=s(x,h)=-a, then $(-1)^{|g|}g^{-1}h=xh^{-1}$ and $hg^{-1}=(-1)^{|h|}h^{-1}x,$; but this is similarly impossible because $|h|\neq|g|.$
@@
Let $d^{\tau}_{z}(g)=\tau(g)gz$ be the central derivation given by the element $z\in A^{1}$. It defines a DG–algebra structure if and only if $\tau(g)\tau(gz)=0\;\forall g\in G$.
@@
defines a DG–algebra structure on $\mathbb{C}[G]$.
@@
It is only necessary to check that the graded Leibniz rule is satisfied. To do this, we can make sure that the given $\tau$ is indeed a group character and then take advantage of Proposition (3.1). It is not difficult to do this by simply going through all the possibilities. ∎
@@
In the graded group algebra $\mathbb{C}[\mathbb{Z}]\cong\bigoplus\mathbb{C}[x^{k}]$, where $\mathbb{Z}=\left\langle{x}\right\rangle$, it is possible to define a DG–algebra structure using the central derivation $d^{\tau}_{x}$, where
@@
Consider the Heisenberg integer group of 3×3 upper–triangular matrices
@@
It is well known that any element of this group can be represented as $y^{b}z^{c}x^{a}$. Then $f(y^{b}z^{c}x^{a})=c\in\mathbb{Z}$ is an epimorphism into a group of integers. From here we get the grading $\mathbb{C}[H_{3}]\cong\bigoplus_{k\in\mathbb{Z}}\mathbb{C}[z^{k}\left\langle{x,y}\right\rangle]$. And the central derivation $d^{\tau}_{z}$, constructed similarly to Example 5.2, gives a DG–algebra structure on $\mathbb{C}[H_{3}]$.